OpenStudy (anonymous):

Choose the replacement for w that makes the equation a true statement. -3w – 24 = -39

OpenStudy (anonymous):

A. 5 B. -5 C. 21 D. -21

OpenStudy (anonymous):

A! -3*5 = -15 -15 + -24 = -39